OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point. 2x2 + xy + 2y2 = 5, (1, 1) (ellipse)

OpenStudy (anonymous):

\[2x^2+xy+2y^2=5\] Implicit differentiation yields \[4x+y+xy'+4y~y'=0\] Solve for \(y'\).

OpenStudy (anonymous):

ok let me try it quick

OpenStudy (anonymous):

i got :

OpenStudy (anonymous):

Not quite: \[4x+y+xy'+4y~y'=0\\ y'~(x+4y)=-(4x+y)\\ y'=-\frac{4x+y}{x+4y}\]

OpenStudy (anonymous):

my apologies- i see where i made my mistake...would this be my final answer

OpenStudy (anonymous):

Nope. \(y'\) gives the slope of the tangent line at any given point. Given the point \((1,1)\), what would \(y'\) be?

OpenStudy (anonymous):

so it would be a - 3/5 slope?